OpenStudy (anonymous):

What is an equation of the line in slopeintercept form? m=2 and the y intercept is (0,3) y=3x+3 y=2x+3 y=2x-3 y=-3x+2

OpenStudy (anonymous):

slope is 2 and given point (x1,y1)=(0,3) the requird eqn is, (y-y1)=m(x-x1) (y-3)=2(x-0) y-3=2x y=2x+3